# Not All Attention is Needed: Parameter and Computation Efficient Tuning for Multi-modal Large Language Models via Effective Attention Skipping

### v2

@@ 5.3.1 Quantitative Analysis
In Tab. 5, we report the results of skipping different numbers of MHAs by EAS. The first observation is that skipping appropriate MHA modules has little impact on performance, e.g. skipping up to 16 MHA modules only has about $1.1\%$ performance drops, strongly suggesting that PIA is a good substitute for MHA. Notably, skipping a certain numbers of MHAs can achieve even better performance than the default LaVIN, e.g. -12 layers with $+2.64$ accuracy. It might be due to the great redundancy of MLLMs on ScienceQA, which also suggests that PIA can help MLLMs learn better patterns from VL data. Again, these results confirm the effectiveness of EAS towards PCETL.
@@ 5.3.2 Qualitative Analysis
Examples of Science Question Answering. In Fig. 4, we visualize the examples of EAS-7B12 and LaVIN-7B for both language-only (left) and image-language questions (right). From these examples, we can see that both EAS-7B and LaVIN-7B correctly answer the question. Meanwhile, they all inherit the strong language ability of LLaMA [50], and can explain the answers fluently. Notably, for the visual question (right), EAS-7B has better and more detailed explanations for the answer, e.g., “The colony’s name comes from the fact that it was founded by people from Hampshire, a county in England”. In contrast, the response of LaVIN is logically incorrect even through its answer is right. Overall, these results confirm again the merits of EAS for the efficient adaption of MLLMs.
6 Limitation
EAS focuses on computation-efficient tuning of MLLMs, thus it has no obvious advantages over PETL methods in parameter efficiency. In this case, EAS is superior in the overall target of PCETL.
@@ 7 Conclusion
In this paper, we propose a novel method for parameter and computation efficient tuning of Multi-modal Large Language Models (MLLMs), named by Effective Attention Skipping. Concretely, we first reveal that not all MHAs are necessary for the efficient adaption of MLLMs, based on which EAS adopts a granular redundancy evaluation scheme. Meanwhile, to avoid the additional computation caused by the adapter-based skip connections, EAS is also equipped with a novel Propagation-of-Information Adapter (PIA), which can not only keep parameter efficiency but also can be re-parameterized into the model without extra latency. To validate EAS, we apply it to a recent MLLM called LaVIN, and conduct extensive experiments on the multi-modal benchmark ScienceQA. The experimental results show that EAS can achieve better performance than LaVIN while speeding up inference by up to 2.18 times.
Appendix A Appendix
A.1 Influence of hidden dimension
Table 6: The impact of different hidden dimensions in PIA for skip-connections of EAS on ScienceQA.
Methods Hidden Dimension Updated Parameters Inference Time Average Performance
EAS-7B8 8 0.04% 1.89s 87.64
16 0.05% 1.91s 87.95
32 0.08% 1.92s 89.79
64 0.22% 1.94s 90.07
Figure 5: Comparison between EAS and DAS on ScienceQA.
Tab. 6 reports the impact of different trainable parameter sizes on EAS via setting the hidden dimension. From the table, we can observe that there is a significant performance increase with more updated parameters, e.g. the performance is improved by $+2.64$ with $0.18\%$ updated parameters more. This result also suggests that replacing redundant MHAs still needs a certain model capacity to accommodate. Even so, the optimal solution, i.e., skipping 12 MHAs, still consumes a very small proportion of parameters, i.e., $0.11\%$, showing high parameter efficiency of EAS.
A.2 Comparison with DAS
We first compare our EAS with the strong baseline DAS under different experimental settings in Fig. 5. The first observation from Fig. 5 is that DAS can skip certain layers of LaVIN with very limited performance drops, e.g., skipping 2-4 layers. Its efficiency gains become more obvious when skipping more than 6 Transformer layers, while the performance also decreases to a certain extent. Compared with DAS, EAS has obvious merits in both performance and efficiency. For instance, with the same number of skipped modules, EAS is consistently faster than DAS by $+44.4\%$ to $51.6\%$. By skipping 8 MHAs, EAS can obtain more obvious performance gains than the default LaVIN, i.e., 89.79 v.s. 89.41, while the inference speed is about $1.92$ times faster. Overall, these results well validate the motivation and designs of EAS towards PCETL of MLLMs.
A.3 The Detailed Results
Table 7: Comparison between EAS and DAS on ScienceQA. The abbreviations for question categories are Natural Science (NAT), Social Science (SOC), Language Science (LAN), Text Context (TXT), Image Context (IMG), No Context (NO), and Grades 1-6 (G1-6), Grades 7-12 (G7-12), respectively. The corner marks of PCETL methods represent the number of skipped modules. The best results of fully supervised and efficient tuning methods are marked in bold and underline, separately, for a fair comparison.
Method Updated Parameters Training Memory Inference Time Subject Context Modality Grade Average
NAT SOC LAN TXT IMG NO G1-6 G7-12
LaVIN-7B [38] 3.8M 32.32G 3.70s 89.25 94.94 85.24 88.51 87.46 88.08 90.16 88.07 89.41
DAS-7B2 [53] 4.2M 32.44G 3.44s 88.68 94.94 86.45 88.03 86.81 88.92 90.20 88.00 89.41
DAS-7B4 [53] 4.6M 31.67G 3.23s 88.99 94.60 85.09 87.88 86.51 88.36 89.72 88.13 89.15
DAS-7B6 [53] 5.0M 30.91G 3.06s 87.30 93.36 82.36 86.12 85.97 85.71 88.18 85.70 87.29
DAS-7B6 [53] 5.4M 30.05G 2.94s 86.50 89.31 84.27 85.63 83.74 86.55 87.41 84.90 86.51
EAS-7B2(Ours) 4.5M 32.82G 2.13s 88.81 95.16 85.36 87.29 86.32 88.99 90.35 87.28 89.25
EAS-7B4(Ours) 5.1M 32.50G 2.08s 88.81 94.83 86.36 87.68 86.81 89.06 90.42 87.67 89.44
EAS-7B6(Ours) 5.7M 32.05G 1.99s 88.99 94.15 85.36 87.88 87.46 88.92 90.68 86.68 89.35
EAS-7B8(Ours) 6.4M 31.63G 1.92s 90.05 94.04 85.82 89.25 87.56 88.78 90.68 88.20 89.79
EAS-7B12(Ours) 7.7M 31.10G 1.69s 89.52 95.61 86.36 88.71 88.25 88.85 91.26 87.67 89.98
We further report the detailed experimental results of DAS and EAS for different numbers of skipped modules. From Tab.7, we observe that the performance of DAS drops faster than the proposed EAS, especially in the TXT benchmark, which states that the skipping of FFN will lead to a weakness in language ability. Meanwhile, we notice that the improvement of inference efficiency in DAS is much slower than in EAS. Overall, the results further prove the effectiveness of our EAS.
Figure 6: Architecture of acceleration models. (a) LaVIN skipped 4 modules by DAS [53]. (b)-(d) LaVIN skipped 4,8,12 MHAs by the proposed EAS. The modules with red index are skipped by the corresponding algorithm.
A.4 The detailed skipping results
We first visualize the accelerated MLLM searched by DAS [53] and our EAS. From Fig. 6, we can observe that the layers skipped by both DAS and EAS tend to be deep layers. While the layers that are considered to have redundancy in EAS are deeper than those in DAS. Furthermore, we can find out that the skipped modules are concentrated behind the middle of MLLM, and gradually expand forward as the number of skipped modules increases.

### v3

@@ 5.3.1 Quantitative Analysis
In Tab. 5.3.1, we report the results of skipping different numbers of MHAs by EAS. The first observation is that skipping appropriate MHA modules has little impact on performance, e.g. skipping up to 16 MHAs only has about $1.1\%$ performance drops, strongly suggesting that PIA is a good substitute for MHA. Notably, skipping a certain number of MHAs can achieve even better performance than the default LaVIN, e.g. -12 layers with +2.64+2.64 accuracy. It might be due to the great redundancy of MLLMs on ScienceQA, which also suggests that PIA can help MLLMs learn better patterns from VL data. We can also observe that EAS can improve efficiency without performance drop within a certain number of skipped layers. Specifically, EAS can skip less than 12 MHAs in LaVIN to accelerate inference latency without performance drop. Again, these results confirm the effectiveness of EAS towards PCETL.
Tab. 8 reports the impact of different trainable parameter sizes on EAS via setting the hidden dimension. From the table, we can observe that more performance gains can be obtained via using more updated parameters, e.g. the performance is improved by +2.64+2.64 with $0.18\%$ updated parameters more. This result also suggests that replacing redundant MHAs still need a certain model capacity to accommodate. Even so, the optimal solution, i.e., skipping 12 MHAs, still consumes a very small proportion of parameters, i.e., $0.11\%$, showing high parameter efficiency of EAS.
@@ 5.3.2 Qualitative Analysis
In Fig. 5, we visualize the language-only and image-language examples of LLaVA-7B with and without our EAS12. From these examples, we can see that via either full tuning or EAS tuning, LLaVA-7B correctly answer the question. Meanwhile, these two settings all keep the strong language ability of LLaMA [touvron2023llama], and can explain the answers fluently. Notably, for the visual question (right), LLaVA-EAS has better and more detailed explanations for the answer, e.g., “The colony’s name comes from the fact that it was founded by people from Hampshire, a county in England”. In contrast, the response of LLaVA-FT is logically incorrect even through its answer is right. Overall, these results confirm again the merits of EAS for the efficient adaption of MLLMs.
@@ 6 Conclusion
In this paper, we propose a novel method for parameter and computation efficient tuning of Multi-modal Large Language Models (MLLMs), named by Effective Attention Skipping. We first reveal that not all MHAs are necessary for the efficient adaption of MLLMs, based on which EAS adopts a granular redundancy evaluation scheme. Meanwhile, to avoid the additional computation caused by the adapter-based skip connections, EAS is also equipped with a novel Propagation-of-Information Adapter (PIA), which can not only keep parameter efficiency but also can be re-parameterized into the model without extra latency. To validate EAS, we apply it to recent MLLMs including LLaVA and LaVIN, and conduct extensive experiments on a set of VL benchmarks and two out-domain datasets. The experimental results show that EAS can achieve better performance than LaVIN while speeding up inference by up to 2.18 times, and EAS also improves the inference speed by 1.20 times while maintaining performance compared to LLaVA.
7 Acknowledgments
This work was supported by the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. U21A20472, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001) and the Fundamental Research Funds for the Central Universities (Xiamen University: No. 20720240053).
8 Declarations
  • Funding: This work was supported by the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. U21A20472, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001) and the Fundamental Research Funds for the Central Universities (Xiamen University: No. 20720240053).
  • Conflict of interest/Competing interests: The authors have no relevant financial or non- financial interests to disclose.
  • Ethics approval and consent to participate: The authors have no relevant ethics approval to disclose.
  • Consent for participate: All authors agreed to participate in this work and made clear contributions.
  • Consent for publication: All authors agreed with the content and that all gave explicit consent to submit and that they obtained consent from the responsible authorities at the institute/organization where the work has been carried out.
  • Data availability: The datasets analysed during the current study are available in these repositories:
    ScienceQA [lu2022learn] https://github.com/lupantech/ScienceQA;
    Slake [conf/isbi/LiuZXMYW21] https://huggingface.co/datasets/BoKelvin/SLAKE;
    AID [journals/tgrs/XiaHHSBZZL17] https://huggingface.co/datasets/blanchon/AID;
    VQAv2 [conf/cvpr/GoyalKSBP17] https://visualqa.org/
    NLVR2 [conf/acl/SuhrZZZBA19] https://lil.nlp.cornell.edu/nlvr/
    Flickr30K [journals/ijcv/PlummerWCCHL17] https://bryanplummer.com/Flickr30kEntities/
  • Code availability: Our code is publicly released at https://github.com/DoubtedSteam/EAS.
  • Author contribution: All authors contributed to the study’s conception and design. Qiong Wu, Weihao Ye and Yiyi Zhou performed the experiment, data collection and analysis. Qiong Wu and Weihao Ye wrote the first draft of the manuscript, and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript. More detailed contributions of each author are listed below: Conceptualization: Qiong Wu, Yiyi Zhou and Xiaoshuai Sun; Methodology: Qiong Wu, Weihao Ye and Yiyi Zhou; Writing-original draft preparation: Qiong Wu and Weihao Ye; Writing-review and editing: Yiyi Zhou, Xiaoshuai Sun and Rongrong Ji; Supervision: Rongrong Ji.